crisulcampo:

There are two pulleys of radius R and r. The distance between their centers is "l". What is the length of the transmission chain?

crisulcampo:

https://postimg.org/image/wli41f7b7/

Falconmaster:

sorry cant open the link computer wont let me

MARC:

http://prntscr.com/dwc715

Falconmaster:

idk sorry not good at that stuff

Falconmaster:

its like 3.14 divide by 2 or somethiing????

crisulcampo:

\[L=\frac{\pi*R}{ 90 }\left[ 180-\cos^{-1} \left( \frac{ R-r }{ l } \right) \right]+\frac{ \pi*r }{ 90 }\cos^{-1} \left( \frac{ R-r }{ l } \right)+2\sqrt{l^{2}-\left( R-r \right)^{2}}\]

crisulcampo:

that's the answer